The pOH of a 0.20 M solution of a weak base is 2.72. What is the Kb and pKb for this base?
Answer #1

Answer I UE table Bars + H2O = BH + oH Initial 0.20M . th Change -n the Equilibrium 0.20 - 2 Kp = 64+ ][64-) - 2? port = 2.12